# Inverse Transfer Multiobjective Optimization

### v1

@@ Abstract.
Transfer optimization enables data-efficient optimization of a target task by leveraging experiential priors from related source tasks. This is especially useful in multiobjective optimization settings where a set of trade-off solutions is sought under tight evaluation budgets. In this paper, we introduce a novel concept of inverse transfer in multiobjective optimization. Inverse transfer stands out by employing probabilistic inverse models to map performance vectors in the objective space to population search distributions in task-specific decision space, facilitating knowledge transfer through objective space unification. Building upon this idea, we introduce the first Inverse Transfer Evolutionary Multiobjective Optimizer (invTrEMO). A key highlight of invTrEMO is its ability to harness the common objective functions prevalent in many application areas, even when decision spaces do not precisely align between tasks. This allows invTrEMO to uniquely and effectively utilize information from heterogeneous source tasks as well. Furthermore, invTrEMO yields high-precision inverse models as a significant byproduct, enabling the generation of tailored solutions on-demand based on user preferences. Empirical studies on multi- and many-objective benchmark problems, as well as a practical case study, showcase the faster convergence rate and modelling accuracy of the invTrEMO relative to state-of-the-art evolutionary and Bayesian optimization algorithms. The source code of the invTrEMO is made available at https://github.com/LiuJ-2023/invTrEMO.
@@
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8
@@ 1. Introduction
In this paper, we introduce a new concept of inverse transfer multiobjective optimization through objective space unification. In conjunction, a first Inverse Transfer Evolutionary Multiobjective Optimizer (invTrEMO) is proposed. The algorithm centres around the construction of probabilistic inverse models that map performance vectors from the objective space to search distributions in task-specific decision space (Cheng et al., 2015; Lin et al., 2022), from which populations of candidate solutions are sampled. The model is iteratively and preferentially updated during the course of a target optimization run with the nondominated solutions found so far, thereby increasing the probability of sampling a higher-quality population in subsequent iterations. The invTrEMO stands out from other transfer optimizers by conducting knowledge transfers through its inverse model. The rationale behind this idea comes from the observation that objective functions of interest frequently coincide in MOPs belonging to the same application area (Tan et al., 2023). Examples abound in science and engineering design applications, among others, where although design variables may vary from one task to the next (e.g., new variables/features are added to a design and obsolete ones removed), objective functions such as overall performance efficiency, cost, robustness or design safety, recur across tasks. The methodological novelty of the invTrEMO lies in treating this common objective space as the unifying bridge, establishing a connection between heterogeneous source-target task pairs as well. In our implementation, this connection is established through the use of principled inverse Transfer Gaussian process (invTGP) models.
@@
  • We present a novel inverse transfer paradigm in multiobjective optimization that melds probabilistic inverse modelling (via invTGPs in our implementation) with evolutionary variation and selection. A key highlight of the paradigm is its ability to facilitate knowledge transfers even across heterogeneous source-target task pairs (with non-identical decision spaces), by harnessing the unification provided by common objective spaces.
@@
[…]
@@ 2. Preliminaries
This section provides an overview of related work and essential background concepts in multiobjective optimization, decomposition-based optimization methods, Gaussian process (GP) modeling and inverse machine learning.
@@ 2.2.4. Inverse Gaussian Processes in MOPs
A simple proof is provided in the Appendix. The proposition establishes that when $\eta=0$, a nondominated solution minimizes the augmented Tchebycheff scalarized function with the preference vector calculated as per (6). This proposition remains approximately valid even for very small values of $\eta$ as in our implementation.
@@ 3.2.1. The invTGP Model
The invTGP model consists of several trainable (hyper-)parameters, including $\sigma_{S,q}$, $\sigma_{T,q}$, $\lambda_{q}$ and the parameters of $k_{q}(\cdot,\cdot)$. These are typically estimated by training the invTGP model using the joint distribution of the source and target tasks (Bonilla et al., 2007). However, this approach can bias the model towards the source rather than the target, especially when $N_{T}^{nd}\ll N_{S}^{nd}$ at the initial stages of the target optimization. To address this issue, we propose a two-step training process that reflects our primary interest in the inverse model’s performance in addressing the target task. In the first step, we learn $\sigma_{T,q}$ and the parameters of $k_{q}(\cdot,\cdot)$ based on $\mathcal{D}_{T,q}^{inv}$ alone, by maximizing the following log marginal likelihood function:
(10) $-\frac{1}{2}\textbf{X}_{T,q}^{\intercal}({\textbf{K}}_{TT,q}+$ $\sigma_{T,q}^{2}\textbf{I}_{N_{T}^{nd}})^{-1}\textbf{X}_{T,q}$
$-\frac{1}{2}\log(|{\textbf{K}}_{TT,q}+\sigma_{T,q}^{2}\textbf{I}_{N_{T}^{nd}}|).$
@@ 3.3. Overall Framework of the invTrEMO
  • Evolutionary Variation and Selection (steps 11-14): Given the selected preference vector $\textbf{w}^{(t)}$ at iteration $t$, an offspring population $\mathcal{U}=\{\textbf{u}^{(l)}\}_{l=1}^{N^{u}}$ is generated by sampling the multivariate Gaussian distribution $[\Psi_{1}^{-1}(\textbf{w}^{(t)}),\Psi_{2}^{-1}(\textbf{w}^{(t)}),\ldots,\Psi_{d_{T}}^{-1}(\textbf{w}^{(t)})]$ in decision space, formed by querying the inverse models with $\textbf{w}^{(t)}$. Thereafter, the upper confidence bound (UCB) (Srinivas et al., 2012) acquisition function is employed as the figure of merit to select one solution $\textbf{x}_{T}^{(t)}$ from $\mathcal{U}$. For minimization problems, the UCB is defined as $-{\mu}_{gp}(\textbf{x}^{(*)})+\beta\cdot{\sigma}_{gp}(\textbf{x}^{(*)})$, where $\beta$ is an appropriate constant. The selected $\textbf{x}_{T}^{(t)}$ is evaluated using the true objective functions, following which $\textbf{x}_{T}^{(t)}$ and $\textbf{f}_{T}(\textbf{x}_{T}^{(t)})$ are appended to $\mathcal{D}_{T}^{M}$.
@@ 4.1. Multiobjective Benchmark Problems
[…]
Figure 1. Influence of $\delta_{1}$ and $\delta_{2}$ on source-target correlations. (a) Pearson correlation coefficients between mDTLZ2-$(1,0)$ and mDTLZ2-$(\delta_{1},0)$ when $\delta_{1}$ is set to 0.3, 0.7, or 0.9. (b) Pearson correlation coefficients between mDTLZ2-$(1,0)$ and mDTLZ2-$(1,\delta_{2})$ when $\delta_{2}$ is set to 0.05, 0.25, or 0.4.
@@
[…]
@@
The choice of $\delta_{1}$ and $\delta_{2}$ values can influence the correlation between the source and target tasks. To approximately quantify this correlation, we compute the Pearson correlation coefficient between the scalarized objective functions of the source and the target task. Fig. 1 illustrates the influence of these two control parameters on mDTLZ2-($\delta_{1}$,$\delta_{2}$). Notably, when $\delta_{1}$ or $\delta_{2}$ deviate significantly from 1 or 0, the source-target correlation is low; conversely, when they are close to 1 or 0, a high source-target correlation is observed. Based on these observations, for each target task, we create source tasks with three levels of source-target correlation by setting $\delta_{1}$ and $\delta_{2}$ to different values. These levels are as follows: highly-correlation source (HS: $\delta_{1}=0.9$ and $\delta_{2}=0.05$), medium-correlation source (MS: $\delta_{1}=0.7$ and $\delta_{2}=0.25$), and low-correlation source (LS: $\delta_{1}=0.3$ and $\delta_{2}=0.4$). The Pearson correlation coefficients for all source-target pairs are summarized in Table 1. To assess invTrEMO’s performance in coping with heterogeneous source-target pairs, we set $d_{S}=6$; thus $d_{S}<d_{T}$. We assume that the first six decision variables of the target task overlap with the variables of the source task. The two remaining variables in the target are non-overlapping.
@@ 4.4. Comparison with State-of-the-Art Evolutionary and Bayesian Optimizers
The above forms a representative set of proficient algorithms in evolutionary and Bayesian optimization, for problems with objectives that are costly to evaluate. Source codes of the invTrEMO and ParEGO-UCB can be found at https://github.com/LiuJ-2023/invTrEMO, with the performance of the ParEGO-UCB favorably verified against the original ParEGO. The source code of PSL-MOBO can be found at https://github.com/Xi-L/PSL-MOBO. The qNEHVI was implemented using BoTorch (Balandat et al., 2020). The implementations of MOEA/D-EGO, K-RVEA, and CSEA were obtained from PlatEMO (Tian et al., 2017). We have summarized the averaged IGD values achieved by the six algorithms after 25, 50, 75, and 100 evaluations (which include the initial $N^{init}$ samples) in Table 2, and the IGD convergence trends are depicted in Fig. 2.
@@
[…]

### v4

# Bayesian Inverse Transfer in Evolutionary Multiobjective Optimization

@@ Abstract.
Transfer optimization enables data-efficient optimization of a target task by leveraging experiential priors from related source tasks. This is especially useful in multiobjective optimization settings where a set of trade-off solutions is sought under tight evaluation budgets. In this paper, we introduce a novel concept of inverse transfer in multiobjective optimization. Inverse transfer stands out by employing Bayesian inverse Gaussian process models to map performance vectors in the objective space to population search distributions in task-specific decision space, facilitating knowledge transfer through objective space unification. Building upon this idea, we introduce the first Inverse Transfer Evolutionary Multiobjective Optimizer (invTrEMO). A key highlight of invTrEMO is its ability to harness the common objective functions prevalent in many application areas, even when decision spaces do not precisely align between tasks. This allows invTrEMO to uniquely and effectively utilize information from heterogeneous source tasks as well. Furthermore, invTrEMO yields high-precision inverse models as a significant byproduct, enabling the generation of tailored solutions on-demand based on user preferences. Empirical studies on multi- and many-objective benchmark problems, as well as a practical case study, showcase the faster convergence rate and modelling accuracy of the invTrEMO relative to state-of-the-art evolutionary and Bayesian optimization algorithms. The source code of the invTrEMO is made available at https://github.com/LiuJ-2023/invTrEMO.
@@
This research is supported partially by Distributed Smart Value Chain programme which is funded under the Singapore RIE2025 Manufacturing, Trade and Connectivity (MTC) Industry Alignment Fund-Pre-Positioning (Award No: M23L4a0001); partially by the Ramanujan Fellowship from the Science and Engineering Research Board, Government of India (Grant No. RJF/2022/000115 ); partially by the Center for Frontier AI research center at the Institute of High Performance Computing, A*STAR.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 1publicationmonth: 8ccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Continuous space search
@@ 1. Introduction
In this paper, we introduce a new concept of inverse transfer multiobjective optimization through objective space unification. In conjunction, a first Inverse Transfer Evolutionary Multiobjective Optimizer (invTrEMO) is proposed for MOPs with expensive black-box objectives. The algorithm centres around the construction of Bayesian inverse Gaussian process (GP) models that map performance vectors from the objective space to search distributions in task-specific decision space (Cheng et al., 2015; Lin et al., 2022), from which populations of candidate solutions are sampled. The model is iteratively and preferentially updated during the course of a target optimization run with the nondominated solutions found so far, thereby increasing the probability of sampling a higher-quality population in subsequent iterations. The invTrEMO stands out from other transfer optimizers by conducting knowledge transfers through its inverse model. A key motivation behind this idea is to address the challenge of heterogeneous decision spaces. In practice, it is possible that decision spaces may not always align across different optimization tasks (Fan et al., 2023). For instance, datasets may originate from experiments exploring various robot skills, each characterized by distinct, potentially overlapping sets of control parameters (Wang et al., 2021). Likewise, in the realm of machine learning model tuning, data sourced from a commercial Bayesian optimization tool (Golovin et al., 2017; Balandat et al., 2020) may exhibit variability due to differences in the hyperparameters tuned by different users and the varied terminologies employed. Notwithstanding such differences in decision space, it is worth noting that the objective functions of interest often coincide in MOPs within the same application domain (Tan et al., 2023). Examples abound in science and engineering design applications, among others, where although design variables may vary from one task to the next (e.g., new variables/features are added to a design and obsolete ones removed), objective functions such as overall performance efficiency, cost, robustness or design safety, recur across tasks. Building upon this observation, the invTrEMO treats this common objective space as a unifying bridge, enabling a connection between heterogeneous source-target task pairs as well. In our implementation, this connection is established through the use of principled inverse Transfer GP (invTGP) models.
@@
  • We present a novel inverse transfer paradigm in multiobjective optimization that melds probabilistic inverse modelling (via invTGPs in our implementation) with evolutionary variation and selection. A key highlight of the paradigm is its ability to facilitate knowledge transfers even across heterogeneous source-target task pairs (with non-identical decision spaces), by harnessing the unification provided by common objective spaces. To the best of our knowledge, this is the first technology to achieve heterogeneous transfer in the context of MOPs with expensive black-box objectives.
@@
  • The effectiveness of our approach is assessed using multi- and many-objective benchmarks from the DTLZ suite (Deb et al., 2005), the DTLZ-1 suite (Ishibuchi et al., 2016), and practical vehicle crashworthiness design examples (Liao et al., 2008). To demonstrate data-efficiency gains, we consider optimization under limited evaluation budgets. Comparisons against state-of-the-art evolutionary and Bayesian optimization algorithms, that are specifically crafted for problems with expensive black-box objectives, confirm invTrEMO’s capacity to boost convergence rates and produce high-precision inverse models in the process.
@@ 2. Preliminaries
This section provides an overview of related work and essential background concepts in multiobjective optimization, decomposition-based optimization methods, GP modeling and inverse machine learning in MOPs.
@@ 2.2.4. Inverse Gaussian Processes in MOPs
A simple proof is provided in the Appendix. Proposition 1 establishes that when $\eta=0$, a nondominated solution minimizes the augmented Tchebycheff scalarized function with the preference vector calculated as per (6). This proposition remains approximately valid even for very small values of $\eta$ as in our implementation.
@@ 3.2.1. The invTGP Model
The invTGP model consists of several trainable (hyper-)parameters, including $\sigma_{S,q}$, $\sigma_{T,q}$, $\lambda_{q}$ and the parameters of $k_{q}(\cdot,\cdot)$. These are typically estimated by training the invTGP model using the joint distribution of the source and target tasks (Bonilla et al., 2007). However, this approach can bias the model towards the source rather than the target, especially when $N_{T}^{nd}\ll N_{S}^{nd}$ at the initial stages of the target optimization. To address this issue, we propose a two step training process that reflects our primary interest in the inverse model’s performance in addressing the target task. In the first step, we learn $\sigma_{T,q}$ and the parameters of $k_{q}(\cdot,\cdot)$ based on $\mathcal{D}_{T,q}^{inv}$ alone, by maximizing the following log marginal likelihood function:
(10) $-\frac{1}{2}\textbf{X}_{T,q}^{\intercal}({\textbf{K}}_{TT,q}+\sigma_{T,q}^{2}\textbf{I}_{N_{T}^{nd}})^{-1}\textbf{X}_{T,q}-\frac{1}{2}\log(|{\textbf{K}}_{TT,q}+\sigma_{T,q}^{2}\textbf{I}_{N_{T}^{nd}}|).$
@@ 3.3. Overall Framework of the invTrEMO
  • Evolutionary Variation and Selection (steps 11-14): Given the selected preference vector $\textbf{w}^{(t)}$ at iteration $t$, an offspring population $\mathcal{U}=\{\textbf{u}^{(l)}\}_{l=1}^{N^{u}}$ is generated by sampling the multivariate Gaussian distribution $[\Psi_{1}^{-1}(\textbf{w}^{(t)}),\Psi_{2}^{-1}(\textbf{w}^{(t)}),\ldots,\Psi_{d_{T}}^{-1}(\textbf{w}^{(t)})]$ in decision space, formed by querying the inverse models with $\textbf{w}^{(t)}$. Thereafter, the upper confidence bound (UCB) (Srinivas et al., 2012) acquisition function is employed as the figure of merit to select [^1] In this procedure, $\textbf{x}_{T}^{(t)}$ is selected by utilizing the predictions provided by the forward GP model. Given that the evaluation cost associated with this predictive process is substantially lower than that of the true expensive evaluations, we operate under the assumption that this process does not incur any significant computational overhead. one solution $\textbf{x}_{T}^{(t)}$ from $\mathcal{U}$. For minimization problems, the UCB is defined as $-{\mu}_{gp}(\textbf{x}^{(*)})+\beta\cdot{\sigma}_{gp}(\textbf{x}^{(*)})$, where $\beta$ is an appropriate constant. The selected $\textbf{x}_{T}^{(t)}$ is evaluated using the true objective functions, following which $\textbf{x}_{T}^{(t)}$ and $\textbf{f}_{T}(\textbf{x}_{T}^{(t)})$ are appended to $\mathcal{D}_{T}^{M}$.
@@
3.4. Summarizing the Transfer Mechanism of invTrEMO
Figure 1. An illustration of the assumed relationship between the source and the target tasks in our problem setup. The common objective space provides the unification through which knowledge transfer could take place, in the inverse setting, between the subset of variables that overlap. Decision variables in distinct tasks are said to overlap if they bear the same physical interpretation in the application domain of interest.
As illustrated in Fig. 1, during the forward mapping, the objective functions of both the source and target tasks map solutions from the decision space to a common objective space. Conversely, the inverse mapping translates points on the PF of the common objective space back to the target task’s decision space given a preference vector as input. In invTrEMO, we approximate this inverse mapping using $Q$ independent invTGPs and $(d_{T}-Q)$ vanilla inverse GPs. Knowledge transfer occurs through the invTGPs, across the $Q$ decision variables that overlap between the source and the target tasks. Unlike a very recent inverse model-based algorithm PSL-MOBO (Lin et al., 2022), these invTGPs are trained utilizing both the source dataset $\mathcal{D}_{S,q}^{inv}$ and the target dataset $\mathcal{D}_{T,q}^{inv}$, thereby encapsulating knowledge from both domains.
In comparison to conventional forward surrogate model-based optimization, knowledge transfer is expected to quickly guide the target search toward performant solutions. The idea is to harness data from source tasks to curtail the need for excessive and expensive exploration of the target search space. This point is experimentally analyzed in Section S-II.1 of the supplementary file, where we show that the offspring sampled from the inverse models’ predicted distribution have significantly lower UCB values than the maximum (as estimated by the forward GP). The invTGP can therefore be said to exploit the source data to achieve more focused target exploration. The effectiveness of the resulting search behavior of the invTrEMO is demonstrated in the experimental studies that follow.
@@ 4.1. Multiobjective Benchmark Problems
We first create eight base problems, each with three objective functions, by modifying the well-known DTLZ and DTLZ-1 benchmark suites (Deb et al., 2005; Ishibuchi et al., 2016). These eight base problems are denoted as mDTLZ1-($\delta_{1}$,$\delta_{2}$), mDTLZ2-($\delta_{1}$,$\delta_{2}$), mDTLZ3-($\delta_{1}$,$\delta_{2}$), mDTLZ4-($\delta_{1}$,$\delta_{2}$), mDTLZ1-1-($\delta_{1}$,$\delta_{2}$), mDTLZ2-1-($\delta_{1}$,$\delta_{2}$), mDTLZ3-1-($\delta_{1}$,$\delta_{2}$), and mDTLZ4-1-($\delta_{1}$,$\delta_{2}$), where $\delta_{1}$ and $\delta_{2}$ serve as two control parameters to vary the problem. The formulations of these eight benchmarks can be found in the supplementary file. By setting $\delta_{1}$ and $\delta_{2}$ to 1 and 0, respectively, we generate four target tasks with $d_{T}=8$, namely mDTLZ1-($1$,00), mDTLZ2-($1$,00), mDTLZ3-($1$,00), mDTLZ4-($1$,00), mDTLZ1-1-($1$,00), mDTLZ2-1-($1$,00), mDTLZ3-1-($1$,00), and mDTLZ4-1-($1$,00). Additionally, a series of source tasks are created by adjusting the values of $\delta_{1}$ and $\delta_{2}$.
Figure 2. Influence of $\delta_{1}$ and $\delta_{2}$ on source-target correlations. (a) Pearson correlation coefficients between mDTLZ2-1-$(1,0)$ and mDTLZ2-1-$(\delta_{1},0)$ when $\delta_{1}$ is set to 0.3, 0.7, or 0.9. (b) Pearson correlation coefficients between mDTLZ2-1-$(1,0)$ and mDTLZ2-1-$(1,\delta_{2})$ when $\delta_{2}$ is set to 0.05, 0.25, or 0.4.
@@
Source Tasks Target Tasks mDTLZ1-1-($1,0$) mDTLZ2-1-($1,0$) mDTLZ3-1-($1,0$) mDTLZ4-1-($1,0$)
@@
The choice of $\delta_{1}$ and $\delta_{2}$ values can influence the correlation between the source and target tasks. To approximately quantify this correlation, we compute the Pearson correlation coefficient between the scalarized objective functions of the source and the target task. Fig. 2 illustrates the influence of these two control parameters on mDTLZ2-1-($\delta_{1}$,$\delta_{2}$). Notably, when $\delta_{1}$ or $\delta_{2}$ deviate significantly from 1 or 0, the source-target correlation is low; conversely, when they are close to 1 or 0, a high source-target correlation is observed. Based on these observations, for each target task, we create source tasks with three levels of source-target correlation by setting $\delta_{1}$ and $\delta_{2}$ to different values. These levels are as follows: highly-correlation source (HS: $\delta_{1}=0.9$ and $\delta_{2}=0.05$), medium-correlation source (MS: $\delta_{1}=0.7$ and $\delta_{2}=0.25$), and low-correlation source (LS: $\delta_{1}=0.3$ and $\delta_{2}=0.4$). The Pearson correlation coefficients for all source-target pairs are summarized in Table 1. To assess invTrEMO’s performance in coping with heterogeneous source-target pairs, we set $d_{S}=6$; thus $d_{S}<d_{T}$. We assume that the first six decision variables of the target task overlap with the variables of the source task. The two remaining variables in the target are non-overlapping.
@@ 4.4. Comparison with State-of-the-Art Evolutionary and Bayesian Optimizers
The above forms a representative set of proficient algorithms in evolutionary and Bayesian optimization, for problems with objectives that are costly to evaluate. Source codes of the invTrEMO and ParEGO-UCB can be found at https://github.com/LiuJ-2023/invTrEMO, with the performance of the ParEGO-UCB favorably verified against the original ParEGO. The source code of PSL-MOBO can be found at https://github.com/Xi-L/PSL-MOBO. The qNEHVI was implemented using BoTorch (Balandat et al., 2020). To ensure fairness in our comparative analysis, the hyperparameters of the forward GPs employed in qNEHVI, ParEGO-UCB, and invTrEMO are configured identically. The implementations of MOEA/D-EGO, K-RVEA, and CSEA were obtained from PlatEMO (Tian et al., 2017). We have summarized the averaged IGD values achieved by the six algorithms after 25, 50, 75, and 100 evaluations (which include the initial $N^{init}$ samples) in Table 2, and the IGD convergence trends are depicted in Fig. 3.
@@
Table 2 indicates that the invTrEMO achieves faster convergence compared to its competitors on a majority of the benchmark problems. The table shows results of the Kruskal-Wallis test at a 0.05 significance level to confirm the statistical significance of the outcomes, where “$+$”, “--”, and $\approx$” denote that the invTrEMO performs better than, worse than, or similar to a competitor. It is noteworthy that the invTrEMO significantly outperforms ParEGO-UCB, MOEA/D-EGO, K-RVEA, and CSEA across all eight target MOPs, and beats qNEHVI and PSL-MOBO on five target problems. qNEHVI and the recent PSL-MOBO only outperform invTrEMO on the last three benchmark problems with inverted triangular PF shapes, i.e., mDTLZ2-1-($1,0$), mDTLZ3-1-($1,0$), and mDTLZ4-1-($1,0$). The above results support the claim that our inverse transfer mechanism can enhance convergence behavior on a wide variety of problems. On problems with inverted PFs, the predefined static preference vectors may not be ideal, and hence the performance of invTrEMO could be further improved in the future with more adaptive objective function weighting schemes. This explains the better performance of qNEHVI and PSL-MOBO over the current version of the invTrEMO on a small subset of problems. The convergence trends in Fig. 3 graphically reinforce the conclusions drawn from Table 2. Figs. S-5 to S-8 of the supplementary file also show the points in objective space corresponding to the nondominated solutions returned by ParEGO-UCB, MOEA/D-EGO, K-RVEA, CSEA, qNEHVI, PSL-MOBO, and the invTrEMO on mDTLZ1-($1,0$) to mDTLZ4-($1,0$) after 100 evaluations. Compared to its competitors, the invTrEMO is found to return solutions that offer better convergence and PF coverage on these benchmark problems.
Figure 3. Comparison of IGD convergence trends averaged over 20 independent runs of ParEGO-UCB, MOEA/D-EGO, K-RVEA, CSEA, qNEHVI, PSL-MOBO, and the invTrEMO. Shaded areas represent one standard deviation on either side of the mean. (a) mDTLZ1-($1,0$). (b) mDTLZ2-($1,0$). (c) mDTLZ3-($1,0$). (d) mDTLZ4-($1,0$). (e) mDTLZ1-1-($1,0$). (f) mDTLZ2-1-($1,0$). (g) mDTLZ3-1-($1,0$). (h) mDTLZ4-1-($1,0$).